OpenStudy (anonymous):

Find all solutions of the equation: sec 4x = 2

OpenStudy (anonymous):

x=1/2

OpenStudy (bahrom7893):

wait is it.. i dont think so.. kumar's wrong

OpenStudy (mertsj):

cos4x=1/2 4x=60 x=15

OpenStudy (mertsj):

If you need all the solutions in a given interval, we had better revisit the solutions.

OpenStudy (anonymous):

The interval is between (o, 2pi]

OpenStudy (mertsj):

Otherwise all the solutions are 4x= 60, 300, 420, 660, 780,1020, 1140,1380 and x= 15, 75, 105, 165,195,255,285,345+2 pi n